\begin{thebibliography}{99}

\bibitem{filomat} R. M. Ali, N. E. Cho, N. Jain and V. Ravichandran, Radii of starlikeness and convexity of functions defined by subordination with fixed second coefficients, Filomat, {\bf 26} (2012), no.~3, 553–-561.

\bibitem{ravi} R. M. Ali, S. K. Lee, V. Ravichandran\ and\ S. Supramaniam, Cofficient estimates for bi-univalent function Ma-Minda starlike and convex functions,  Appl. Math. Lett., {\bf 25} (2012), 344--351.

 \bibitem{ali11} R. M. Ali, S. Nagpal\ and\ V. Ravichandran, Second-order differential subordination for analytic functions with fixed initial coefficient, Bull. Malays. Math. Sci. Soc. (2) {\bf 34} (2011), no.~3, 611--629.

\bibitem{branan} A. Brannan and J. G. Clunie, Aspects of contemporary complex analysis Proceedings of the NATO Advanced Study Institute held at the University of Durham, Durham, July 1–20, 1979, Academic Press New York, London, 1980.

 \bibitem{branan0} D. A. Brannan\ and\ T. S. Taha, On some classes of bi-univalent functions, in {\it Mathematical analysis and its applications (Kuwait, 1985)}, 53--60, KFAS Proc. Ser., 3 Pergamon, Oxford.

\bibitem{branan1} D. A. Brannan\ and\ T. S. Taha, On some classes of bi-univalent functions, Studia Univ. Babe\c s-Bolyai Math. {\bf 31} (1986), no.~2, 70--77.
\bibitem{bhs} D. Bshouty, W. Hengartner\ and\ G. Schober, Estimates for the Koebe constant and the second coefficient for some classes of univalent functions, Canad. J. Math. {\bf 32} (1980), no.~6, 1311--1324.


\bibitem{chen} Sh. Chen, S. Ponnusamy\ and\ X. Wang, Coefficient estimates and Landau-Bloch's constant for planar harmonic mappings, Bull. Malays. Math. Sci. Soc. (2) {\bf 34} (2011), no.~2, 255--265.

\bibitem{cho10} N. E. Cho\ and\ O. S. Kwon, A class of integral operators preserving subordination and superordination, Bull. Malays. Math. Sci. Soc. (2) {\bf 33} (2010), no.~3, 429--437.


\bibitem{frasin}B. A. Frasin\ and\ M. K. Aouf, New subclasses of bi-univalent functions, Appl. Math. Lett. {\bf 24} (2011), no.~9, 1569--1573.

\bibitem{good}A. W. Goodman, An invitation to the study of univalent and multivalent functions, Internat. J. Math. Math. Sci. {\bf 2} (1979), no.~2, 163--186.
\bibitem{jano} W. Janowski, Extremal problems for a family of functions with positive real part and for some related families, Ann. Polon. Math. {\bf 23} (1970/1971), 159--177.
\bibitem{ked89} A. W. Kedzierawski, Some remarks on bi-univalent functions, Ann. Univ. Mariae Curie-Sk\l odowska Sect. A {\bf 39} (1985), 77--81 (1988).

\bibitem{ked88} A. W. Kedzierawski\ and\ J. Waniurski, Bi-univalent polynomials of small degree, Complex Variables Theory Appl. {\bf 10} (1988), no.~2-3, 97--100.
\bibitem{koef69} F. R. Keogh and E. P. Merkes, A coefficient inequality for certain classes of analytic functions, Proc. Amer. Math. Soc. {\bf 20} (1969), 8--12.
\bibitem{lewin} M. Lewin, On a coefficient problem for bi-univalent functions, Proc. Amer. Math. Soc. {\bf 18} (1967), 63--68.

\bibitem{liu11} J.-L. Liu, Certain sufficient conditions for strongly starlike functions associated with an integral operator, Bull. Malays. Math. Sci. Soc. (2) {\bf 34} (2011), no.~1, 21--30.
\bibitem{minda}  W. C. Ma\ and\ D. Minda, A unified treatment of some special classes of univalent functions, in {\it Proceedings of the Conference on Complex Analysis (Tianjin, 1992)}, 157--169, Conf. Proc. Lecture Notes Anal., I Int. Press, Cambridge, MA.
\bibitem{neta} E. Netanyahu, The minimal distance of the image boundary from the origin and the second coefficient of a univalent function in $|z|<\,1$, Arch.
     Rational Mech. Anal. {\bf 32} (1969), 100--112.
%\bibitem{pomer} C. Pommerenke, {\it Univalent functions}, Vandenhoeck \&\ Ruprecht, G\"ottingen, 1975.

\bibitem{ravi05} V. Ravichandran,\ Y. Polatoglu,\ M. Bolcal,\ A. Sen, Certain subclasses of starlike and convex functions of complex order, Hacet. J. Math. Stat. {\bf 34} (2005), 9--15.

\bibitem{rob} M. I. S. Robertson, On the theory of univalent functions, Ann. of Math. (2) {\bf 37} (1936), no.~2, 374--408.


\bibitem{smith1} H. V. Smith, Bi-univalent polynomials, Simon Stevin {\bf 50} (1976/77), no.~2, 115--122.

\bibitem{smith2}    H. V. Smith, Some results/open questions in the theory of bi-univalent functions, J. Inst. Math. Comput. Sci. Math. Ser. {\bf 7} (1994), no.~3, 185--195.

\bibitem{srivastava} H. M. Srivastava, A. K. Mishra\ and\ P. Gochhayat, Certain subclasses of analytic and bi-univalent functions, Appl. Math. Lett. {\bf 23} (2010), no.~10, 1188--1192.

\bibitem{sty81} D. Styer\ and\ J. Wright, Result on bi-univalent functions, Proc. Amer. Math. Soc., Vol.82, No~2, 1981, 243--248.
\bibitem{suff69} T. J. Suffridge, A coefficient problem for a class of univalent functions,
Michigan Math. J. 16(1969), 33-42.
\bibitem{samni}S. Supramaniam, R.M. Ali, S. K. Lee and V. Ravichandran, Convolution and differential subordination for multivalent functions, Bull. Malays. Math. Sci. Soc. (2) {\bf 32} (2009), no.~3, 351--360.

%
%\bibitem{taha} T. S. Taha, Topics in univalent function theory, Ph.D thesis, Univ. of London, 1981.
\bibitem{tan}D. L. Tan, Coefficient estimates for bi-univalent functions, Chinese Ann. Math. Ser. A {\bf 5} (1984), no.~5, 559--568.

 \bibitem{xu}Q.-H. Xu, Y.-C. Gui and H. M. Srivastava, Coefficient estimates for a certain subclass of analytic and bi-univalent functions,   Appl. Math. Lett. {\bf 25} (2012), 990--994.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{cite,graphics,mathptmx}
\newcommand{\set}[1]{\left\{#1\right\}}

\numberwithin{equation}{section}

\newtheorem{theorem}{Theorem}[section]

\theoremstyle{remark}
 \newtheorem{remark}[theorem]{Remark}

\newtheorem{definition}[theorem]{Definition}

\begin{document}
 
\title{Estimates for the Initial Coefficients of Bi-univalent Functions}
\author[S. Sivaprasad Kumar]{S. Sivaprasad Kumar}

\address{Department of Applied Mathematics,
Delhi Technological University, Delhi---110042, India}
\email{spkumar@dce.ac.in}

\author[V. Kumar]{Virendra Kumar}
\address{Department of Applied Mathematics,
Delhi Technological University, Delhi---110042, India}
\email{vktmaths@yahoo.in}

\author{V. Ravichandran}
\address{Department of Mathematics, University of Delhi, Delhi---110007, India}
\email{vravi@maths.du.ac.in}

\begin{abstract}A bi-univalent function is a univalent function defined on the  unit disk with its inverse also univalent on the unit disk.  In the present investigation, estimates for the initial coefficients are obtained  for   bi-univalent functions belonging to certain classes defined by subordination and  relevant connections with earlier results are pointed out.
\end{abstract}

\keywords{Univalent functions, bi-univalent functions, coefficient estimate, subordination.}

\subjclass[2010]{30C45, 30C80}

\maketitle

\section{Introduction}
Let $\mathcal{A}$ be the class of analytic functions defined  on the open unit disk $\mathbb{D}:=\{z\in\mathbb{C}:|z|<1\}$ and normalized by the conditions $f(0)=0$ and $f'(0)=1$. A function $f\in \mathcal{A}$  has Taylor's series expansion of the form \begin{equation}\label{eq1}
f(z)=z+\sum^{\infty}_{n=2}a_nz^n.\end{equation}
The class of all  univalent functions in the open unit disk $\mathbb{D}$ of the form (\ref{eq1})
 is denoted by $\mathcal{S}$. Determination of the bounds for  the coefficients $a_n$  is an important problem in
 geometric function theory as they give information about the geometric properties of these functions.
  For example, the bound for the second coefficient $a_2$ of  functions in $\mathcal{S}$ gives the growth
  and distortion bounds as well as covering theorems. Some coefficient related problems were investigated
  recently in \cite{filomat,ali11,chen,cho10,liu11,samni}. 

Since univalent functions are one-to-one, they are invertible but their inverse functions need not be defined on the entire unit disk $\mathbb{D}$. In fact, the famous Koebe one-quarter theorem ensures that the image of the unit disk $\mathbb{D}$ under every function $f\in\mathcal{S}$ contains a disk of radius $1/4$. Thus, inverse of every function $f\in\mathcal{S}$ is defined on a disk, which contains the disk $|z|<1/4$. It can also be easily verified that
\begin{equation}\label{ei}F(w):=f^{-1}(w)=w-a_2w^2+(2a_2^2-a_3)w^3-(5a_2^2-5a_2a_3+a_4)w^4+
\cdots\end{equation}
in some disk of radius at least 1/4.  A function $f\in\mathcal{A}$ is called \emph{bi-univalent}  in $\mathbb{D}$ if both $f$ and $f^{-1}$ are univalent in $\mathbb{D}$. In 1967, Lewin \cite{lewin} introduced the class $\sigma$ of bi-univalent analytic functions and showed that the second coefficient of every $f\in\sigma$  satisfy the inequality $|a_2|\leq1.51$. Let $\sigma_1$ be the class of all functions $f=\phi\circ\psi^{-1}$ where $\phi,\psi$ map $\mathbb{D}$ onto a domain containing $ \mathbb{D}$ and $\phi'(0)=\psi'(0)$. In 1969, Suffridge \cite{suff69} gave a function in $\sigma_1\subset\sigma$, satisfying $a_2=4/3$ and conjectured that $|a_2|\leq 4/3$ for all functions in $\sigma$. In 1969, Netanyahu \cite{neta} proved this conjecture for the subclass $\sigma_1$. Later in 1981, Styer and Wright \cite{sty81} disproved the conjecture of Suffridge \cite{suff69} by showing $a_2>4/3$ for some function in $\sigma$. Also see \cite{bhs} for an example to show $\sigma\neq \sigma_1$. For results on bi-univalent polynomial, see \cite{smith1,ked88}. In 1967, Brannan \cite{branan} conjectured that $|a_2|\leq \sqrt{2}$ for $f\in \sigma$. In 1985, Kedzierawski {\cite [Theorem 2]{ked89}} proved this conjecture for a special case when the function  $f $  and $f^{-1}$  are starlike functions. In $1985$, Tan \cite{tan} obtained the bound for $a_2$ namely $|a_2|\leq 1.485$ which is the best known estimate for functions in the class $\sigma$. For some open problems and survey, see \cite{good,smith2}. In $1985$, Kedzierawski \cite{ked89} proved the following:
 $$|a_2|\leq\left\{
   \begin{array}{ll}
     1.5894, & \hbox{$f\in \mathcal{S},\; f^{-1}\in \mathcal{S}$;} \\
     \sqrt2, & \hbox{$f\in \mathcal{S}^*,\; f^{-1}\in \mathcal{S}^*$;} \\
     1.507, & \hbox{$f\in \mathcal{S}^*,\; f^{-1}\in \mathcal{S}$;} \\
     1.224, & \hbox{$f\in \mathcal{K},\; f^{-1}\in \mathcal{S}$,}
   \end{array}
 \right.$$
where $\mathcal{S}^*$ and $\mathcal{K}$ denote the well-known classes of starlike and convex functions in $\mathcal{S}.$

 Let us recall now various definitions required in sequel. An analytic  function $f$ is  \emph{subordinate} to another analytic function $g$, written $f\prec g$, if there is an analytic function $w$ with $|w(z)|\leq |z|$ such that $f =g\circ w$. If $g$ is univalent, then $f\prec g$ if and only if $f(0)=g(0)$ and $f(\mathbb{D})\subseteq g(\mathbb{D})$. Let $\varphi$ be an analytic univalent function in $\mathbb{D}$ with positive real part  and $\varphi(\mathbb{D})$ be symmetric with respect to the real axis, starlike with respect to $\varphi(0)=1$  and $\varphi'(0)>0.$ Ma and Minda \cite{minda} gave a unified presentation of various subclasses of starlike and convex functions by introducing the classes $\mathcal{S}^*(\varphi)$ and $\mathcal{K}(\varphi)$ of functions $f\in \mathcal{S}$ satisfying $zf'(z)/f(z)\prec\varphi(z)$ and $1+zf''(z)/f'(z)\prec\varphi(z)$ respectively, which includes several well-known classes as special case. For example, when $\varphi(z)=(1+A z)/(1+Bz)\; (-1\leq B<A\leq 1),$ the class $\mathcal{S}^*(\varphi)$ reduces to the class $\mathcal{S}^*[A,B]$ introduced by Janowski \cite{jano}.  For $0\leq\beta<1$, the classes $\mathcal{S}^*(\beta):=\mathcal{S}^*((1+(1-2\beta)z)/(1-z))$
 and $\mathcal{K}(\beta):=\mathcal{K}((1+(1-2\beta)z)/(1-z))$ are starlike and convex functions of order $\beta.$
Further let $\mathcal{S}^*:=\mathcal{S}^*(0)$ and $\mathcal{K}:=\mathcal{K}(0)$ are the classes of starlike and convex functions respectively. The class of strongly starlike functions $\mathcal{S}^*_\alpha:=\mathcal{S}^*(((1+z)/(1-z))^\alpha)$ of order $\alpha,\; 0<\alpha\leq1.$ Denote by $\mathcal{R}(\varphi)$ the class of all functions satisfying $f'(z)\prec\varphi(z)$ and let $\mathcal{R}(\beta):=\mathcal{R}((1+(1-2\beta)z)/(1-z))$ and $\mathcal{R}:=\mathcal{R}(0)$.

For $ 0\leq\beta<1$, a function $f\in \sigma$ is in the class $\mathcal{S}^*_{\sigma}(\beta)$ of \emph{bi-starlike  function of order $\beta$}, or $\mathcal{K}_{\sigma}(\beta)$ of\emph{ bi-convex function of order $\beta$} if both $f$ and $f^{-1}$ are respectively starlike or convex functions of order $\beta$. For $0<\alpha\leq 1$,  the function $f\in\sigma$ is   \emph{strongly bi-starlike function of order $\alpha$} if both the functions $f$ and $f^{-1}$ are strongly starlike functions of order $\alpha$. The class of all such functions is denoted by $\mathcal{S}^*_{\sigma,\alpha}$. These classes were introduced by  Brannan and Taha~\cite{branan1} in 1985 (see also \cite{branan0}). They obtained estimates on the initial coefficients $a_2$ and $a_3$ for functions in these classes. Recently, Ali \emph{et al.}\ \cite{ravi} extended the results of   Brannan and Taha \cite{branan1}  by generalizing their classes using subordination. For some related results, see \cite{srivastava,frasin,xu}. For the various applications of subordination one can refer to \cite{filomat,ali11,cho10,liu11,samni} and the references cited therein.

Motivated by  Ali \emph{et al.}\ \cite{ravi} in this paper estimates for the initial coefficient $a_2$ of bi-univalent functions belonging to the  class $\mathcal{R}_\sigma(\lambda, \varphi)$ as well as estimates on $a_2$ and $a_3$ for functions in classes $\mathcal{S^*}_\sigma(\varphi)$ and $K_{\sigma}(\varphi)$, defined later, are obtained. Further work of Kedzierawski \cite{ked89} actuates us to derive the estimates on  initial coefficients $a_2$ and $a_3$ when $f$ is in the some subclass of univalent functions and $f^{-1}$ belongs to some other subclass of univalent functions. Our results generalize several well-known results in \cite{ravi, frasin,ked89,srivastava}, which are pointed out here.

\section{Coefficient estimates}
Throughout this paper, we assume that  $\varphi$ is an analytic function in $\mathbb{D}$ of the form \begin{equation}\label{eb}\varphi(z)=1+B_1z+B_2z^2+B_3z^3+\cdots \;\;\text{with}\;\;B_1>0,\;{\text {and}}\; B_2\;{\text{ is any real number}} .\end{equation}

\begin{definition}Let $\lambda\geq0$.  A function $f\in \sigma$ given by (\ref{eq1}) is  in the class $\mathcal{R}_\sigma(\lambda, \varphi)$,  if it satisfies
\[(1-\lambda)\frac{f(z)}{z}+\lambda f'(z)\prec\varphi(z) \quad \text{ and}\quad  (1-\lambda)\frac{F(w)}{w}+\lambda F'(w)\prec\varphi(w).\] \end{definition}

The class $\mathcal{R}_\sigma(\lambda, \varphi)$ includes many earlier classes, which are mentioned below:
\begin{enumerate}
\item $\mathcal{R}_\sigma(\lambda, (1+(1-2\beta)z)/(1-z))=\mathcal{R}_{\sigma}(\lambda,\beta)\;$  $(\lambda\geq 1;\;0\leq\beta<1)$    \cite[Definition 3.1]{frasin}

\item $\mathcal{R}_\sigma(\lambda, ((1+z)/(1-z))^\alpha)=\mathcal{R}_{\sigma,\alpha}(\lambda)\;\; (\lambda\geq 1;\;0<\alpha\leq1)$    \cite[Definition 2.1]{frasin}

\item $\mathcal{R}_\sigma(1,\varphi)
      =\mathcal{R}_{\sigma}(\varphi)$  \cite[p. 345]{ravi}.
\item $\mathcal{R}_\sigma(1,(1+(1-2\beta)z)/(1-z))
      =\mathcal{R}_\sigma(\beta)\;\; (0\leq\beta<1)$  \cite[Definition 2]{srivastava}
\item  $\mathcal{R}_\sigma(1,((1+z)/(1-z))^\alpha)
      =\mathcal{R}_{\sigma,\alpha}\;\; (0<\alpha\leq1)$  \cite[Definition 1]{srivastava}

\end{enumerate}
 Our first result provides estimate for the coefficient $a_2$ of functions $f\in\mathcal{R}_\sigma(\lambda, \varphi)$.

\begin{theorem}\label{th1} If $f\in\mathcal{R}_\sigma(\lambda, \varphi)$, then
\begin{align}\label{e1}|a_2|&\leq \sqrt\frac{B_1+|B_1-B_2|}{1+2\lambda}
\end{align}
\end{theorem}

\begin{proof}Since $f\in\mathcal{R}_\sigma(\lambda, \varphi)$,   there exist two analytic functions $r, s:\mathbb{D}\rightarrow\mathbb{D}$,  with $r(0)=0=s(0)$,  such that
\begin{equation}\label{p1}(1-\lambda)\frac{f(z)}{z}+\lambda f'(z)=\varphi(r(z))\;\text {and}\; (1-\lambda)\frac{F(w)}{w}+\lambda F'(w)=\varphi(s(z)).\end{equation}
Define the functions $p$ and $q$
 by
 {\small\begin{equation}\label{pq}
    p(z)=\frac{1+r(z)}{1-r(z)}= 1+p_1z+p_2z^2+p_3z^3+\cdots \;\;\text {and}\;\; q(z)=\frac{1+s(z)}{1-s(z)}= 1+q_1z+q_2z^2+q_3z^3+\cdots,
 \end{equation}}
 or equivalently,
 \begin{equation}\label{p2}
  r(z)=\frac{p(z)-1}{p(z)+1}=\frac{1}{2}\left(p_1z+\Big(p_2-\frac{p_1^2}{2}\Big)z^2+
  \Big(p_3+\frac{p_1}{2}\big(\frac{p_1^2}{2}-p_2\big)-\frac{p_1p_2}{2}\Big)z^3+\cdots\right)
 \end{equation}
 and
\begin{equation}\label{p3}
  s(z)=\frac{q(z)-1}{q(z)+1}=\frac{1}{2}\left(q_1z+\Big(q_2-\frac{q_1^2}{2}\Big)z^2+ \Big(q_3+\frac{q_1}{2}\big(\frac{q_1^2}{2}-q_2\big)-\frac{q_1q_2}{2}\Big)z^3+\cdots\right).
 \end{equation}
It is clear that $p$ and $q$ are analytic in $\mathbb{D}$ and $p(0)=1=q(0)$. Also $p$ and $q$  have positive real part in $\mathbb{D}$, and hence $|p_i|\leq2$ and  $|q_i|\leq2$. In the view of (\ref{p1}), (\ref{p2}) and (\ref{p3}), clearly
 \begin{equation}\label{p4}(1-\lambda)\frac{f(z)}{z}+\lambda f'(z)=\varphi\left(\frac{p(z)-1}{p(z)+1}\right)\;\;\text {and}\;\; (1-\lambda)\frac{F(w)}{w}+\lambda F'(w)=\varphi\left(\frac{q(w)-1}{q(w)+1}\right).\end{equation}
On expanding (\ref{eb}) using (\ref{p2}) and (\ref{p3}), it is evident that
\begin{multline}\label{p5}
\varphi\left(\frac{p(z)-1}{p(z)+1}\right)= 1+\frac{1}{2}B_1p_1z+\left(\frac{1}{2}B_1\big(p_2-\frac{1}{2}p_1^2\big)
   +\frac{1}{4}B_2p_1^2\right)z^2+\cdots.
\end{multline}
and
\begin{multline}\label{p6}
\varphi\left(\frac{q(w)-1}{q(w)+1}\right)= 1+\frac{1}{2}B_1q_1w+\left(\frac{1}{2}B_1\big(q_2-\frac{1}{2}q_1^2\big)
   +\frac{1}{4}B_2q_1^2\right)w^2+\cdots.
\end{multline}
Since $f\in\sigma$ has the Maclaurin series given by (\ref{eq1}), a computation shows that its inverse $F=f^{-1}$ has the expansion given by (\ref{ei}). It follows from (\ref{p4}), (\ref{p5}) and (\ref{p6}) that
\[(1+\lambda) a_2=\frac{1}{2}B_1p_1,\]
\begin{equation}\label{eq12}
 (1+2\lambda) a_3 =\frac{1}{2}B_1\Big(p_2-\frac{1}{2}p_1^2\Big)+\frac{1}{4}B_2p_1^2,
\end{equation}
\[ -(1+\lambda) a_2=\frac{1}{2}B_1q_1,\]
\begin{equation}\label{eq16}
 (1+2\lambda) (2a_2^2-a_3) =\frac{1}{2}B_1\left(q_2-\frac{1}{2}q_1^2\right)+\frac{1}{4}B_2q_1^2.
\end{equation}
Now  (\ref{eq12}) and (\ref{eq16}) yield
 \begin{equation}\label{eq17a}
 8(1+2\lambda)a_2^2=2(p_2+q_2)B_1+(B_2-B_1)(p_1^2+q_1^2).
\end{equation}
Finally an application of the known results, $|p_i|\leq2$ and  $|q_i|\leq2$
 in (\ref{eq17a}) yields the desired estimate of $a_2$ given by (\ref{e1}).
\end{proof}
\begin{remark} Let $\varphi(z)=(1+(1-2\beta)z)/(1-z),\; 0\leq\beta<1.$ So $B_1=B_2=2(1-\beta).$
 When $\lambda=1$, Theorem \ref{th1} gives the estimate $|a_2|\leq \sqrt {2(1-\beta)/3}$ for functions in the class $\mathcal{R}_\sigma(\beta)$ which coincides with the result \cite[Corollary 2]{xu} of Xu \emph{et al}.
In particular if $\beta=0$, then above estimate becomes $|a_2|\leq \sqrt{2/3}\approx 0.816$ for functions $f\in \mathcal{R}_\sigma(0).$ Since the estimate on $|a_2|$ for $f\in  \mathcal{R}_\sigma(0) $ is improved over the conjectured estimate   $|a_2|\leq\sqrt{2}\approx 1.414$ for $f\in \sigma$, the functions in $ \mathcal{R}_\sigma(0) $ are not the candidate for the sharpness of the estimate in the class $\sigma$.
\end{remark}
\begin{definition}  A function $f\in \sigma$  is  in the class $\mathcal{S^*}_\sigma(\varphi)$,  if it satisfies
 $$\frac{zf'(z)}{f(z)}\prec\varphi(z)\quad \text{and}\quad \frac{wF'(w)}{F(w)}\prec\varphi(w).$$
 \end{definition}
 Note that for a suitable choice of $\varphi$,  the class $\mathcal{S^*}_\sigma(\varphi)$, reduces to the following well-known classes:
 \begin{enumerate}
   \item $\mathcal{S^*}_\sigma((1+(1-2\beta)z)/(1-z))=\mathcal{S^*}_\sigma(\beta) \quad(0\leq\beta<1).$
   \item $\mathcal{S^*}_\sigma(\left((1+z)/(1-z)\right)^\alpha)
    =\mathcal{S}^*_{\sigma,\alpha} \quad(0<\alpha\leq 1).$
 \end{enumerate}
\begin{theorem} \label{th2}If $f\in\mathcal{S^*}_\sigma(\varphi)$, then
$$|a_2|\leq\min\set{\sqrt{B_1+|B_2-B_1|}, \sqrt{ \frac{B_1^2+B_1+|B_2-B_1|}{2}}, \frac{B_1\sqrt{B_1}}{\sqrt{B_1^2+|B_1-B_2|}}}$$ and
$$|a_3|\leq\min\set{ B_1+|B_2-B_1|, \frac{B_1^2+B_1+|B_2-B_1|}{2}, R},$$ where
$$R:= \frac{1}{4}\left(B_1+3B_1\max\set{1;\left|\frac{B_1-4B_2}{3B_1}\right|}\right).$$
\end{theorem}

\begin{proof} Since $f\in \mathcal{S}^*_\sigma(\varphi)$,  there are analytic functions $r, s:\mathbb{D}\rightarrow\mathbb{D}$,  with $r(0)=0=s(0)$,  such that
\begin{equation}\label{p1.1}\frac{zf'(z)}{f(z)}=\varphi(r(z))\;\text {and}\; \frac{wF'(w)}{F(w)} =\varphi(s(z)).\end{equation}
Let $p$ and $q$ be defined as in (\ref{pq}),
then it is clear from (\ref{p1.1}), (\ref{p2}) and (\ref{p3}) that
\begin{equation}\label{p4.1}\frac{zf'(z)}{f(z)}=\varphi\left(\frac{p(z)-1}{p(z)+1}\right)\;\;\text {and}\;\; \frac{wF'(w)}{F(w)}=\varphi\left(\frac{q(z)-1}{q(z)+1}\right).\end{equation}
It follows from (\ref{p4.1}), (\ref{p5}) and (\ref{p6}) that
\begin{equation}\label{eq11.1}
 a_2=\frac{1}{2}B_1p_1,
\end{equation}
\begin{equation}\label{eq12.1}
 2a_3 =\frac{B_1p_1}{2}a_2+\frac{1}{2}B_1\left(p_2-\frac{1}{2}p_1^2\right)+\frac{1}{4}B_2p_1^2,
\end{equation}
\begin{equation}\label{eq15.1}
 -a_2=\frac{1}{2}B_1q_1
\end{equation}
and
\begin{equation}\label{eq16.1}
 4a_2^2-2a_3=-\frac{B_1q_1}{2}a_2+
 \frac{1}{2}B_1\left(q_2-\frac{1}{2}q_1^2\right)+\frac{1}{4}B_2q_1^2.
\end{equation}
The equations (\ref{eq11.1}) and (\ref{eq15.1}) yield
\begin{equation}\label{eq17.1}
 p_1=-q_1,
\end{equation}
\begin{equation}\label{eq18.10}
 8a_2^2=(p_1^2+q_1^2)B_1^2
\end{equation}
and
\begin{equation}\label{eq18.1}
 2a_2=\frac{B_1(p_1-q_1)}{2}.
\end{equation}
From (\ref{eq12.1}), (\ref{eq16.1}) and (\ref{eq18.1}), it follows that
\begin{equation}\label{eq19}
 8a_2^2=2B_1(p_2+q_2)+(B_2-B_1)(p_1^2+q_1^2).
\end{equation}
Further a computation using (\ref{eq12.1}), (\ref{eq16.1}), (\ref{eq11.1}) and (\ref{eq17.1}) gives
\begin{equation}\label{eq19.1}
 16a_2^2=2B_1^2q_1^2+2B_1(p_2+q_2)+(B_2-B_1)(p_1^2+q_1^2).
\end{equation}
Similarly a computation using (\ref{eq12.1}), (\ref{eq16.1}), (\ref{eq18.1}) and (\ref{eq18.10}) yields
\begin{equation}\label{eq19.10}
 4(B_1^2-B_2+B_1)a_2^2=B_1^3(p_2+q_2).
\end{equation}
Now (\ref{eq19}), (\ref{eq19.1}) and (\ref{eq19.10}) yield the desired estimate on $a_2$ as asserted in the theorem. To find estimate for $a_3$ subtract (\ref{eq12.1}) from (\ref{eq16.1}), to get
\begin{equation}\label{eq19.2}
 -4a_3=-4a_2^2+\frac{B_1(q_2-p_2)}{2}.
\end{equation}
 Now a computation using (\ref{eq19.1}) and (\ref{eq19.2}) leads to
\begin{equation}\label{eq19.31}
 16a_3=2B_1^2q_1^2+4B_2p_2+(B_1-B_2)(p_1^2+q_1^2).
\end{equation}
From (\ref{eq11.1}), (\ref{eq12.1}), (\ref{eq15.1}) and (\ref{eq16.1}), it follows that

\begin{eqnarray}\label{eq19.33}
  4a_3 &=& \frac{B_1}{2}(3p_2+q_2)+(B_2-B_1)p_1^2 \\ \label{eq19.3}
   &=& \frac{B_1q_2}{2}+\frac{3B_1}{2}\left(p_2-\frac{2(B_1-B_2)}{3B_1}p_1^2\right).
\end{eqnarray} On applying the result of Keogh and Merkes \cite{koef69}(see also \cite{ravi05}), that is for any complex number $v$, $|p_2-vp_1^2|\leq 2 \max\{1;|2v-1|\}$, along with $|q_2|\leq2$ in (\ref{eq19.3}), we obtain
\begin{equation}\label{eq19.30}
 4|a_3|\leq B_1+3B_1\max\set{1;\left|\frac{B_1-4B_2}{3B_1}\right|}.
\end{equation}
 Now the desired estimate on $a_3$ follows from (\ref{eq19.31}), (\ref{eq19.33}) and (\ref{eq19.30}) at once.\end{proof}
\begin{remark}  If $f\in\mathcal{S^*}_\sigma(\beta)\;\;(0\leq\beta<1)$, then from Theorem~\ref{th2} it is evident that

{\small\begin{equation}\label{1}
|a_2|\leq\min\set{\sqrt{2(1-\beta)}, \sqrt{(1-\beta)(3-2\beta)}}=\left\{
  \begin{array}{ll}
    \sqrt{2(1-\beta)}, & \hbox{$0\leq\beta\leq1/2$;} \\
    \sqrt{(1-\beta)(3-2\beta)}, & \hbox{$1/2\leq\beta<1$.}
  \end{array}
\right.
\end{equation}}
Recall Brannan and Taha's~\cite[Theorem 3.1]{branan0} coefficient estimate, $|a_2|\leq \sqrt{2(1-\beta)}$ for functions $f\in\mathcal{S^*}_\sigma(\beta),$ who claimed that their estimate is better than the estimate $|a_2|\leq 2(1-\beta)$, given by Robertson~\cite{rob}. But their claim is true only when $0\leq\beta\leq 1/2.$ Also it may noted that our estimate for $a_2$ given in (\ref{1}) improves the estimate given by Brannan and Taha~\cite[Theorem 3.1]{branan0}.

Further if we take $\varphi(z)=((1+z)/(1-z))^\alpha,\;0<\alpha\leq1$ in Theorem~\ref{th2}, we have $B_1=2\alpha$ and $B_2=2\alpha^2$. Then we obtain the estimate on $a_2$ for functions $f\in\mathcal{S^*}_{\sigma,\alpha}$ as:
$$|a_2|\leq\min\set{\sqrt{4\alpha-2\alpha^2}, \sqrt{\alpha^2+2\alpha}, \frac{2\alpha}{\sqrt{1+\alpha}}}=\frac{2\alpha}{\sqrt{1+\alpha}}.$$
 Note that Brannan and Taha~\cite[Theorem 2.1]{branan0} gave the same estimate $|a_2|\leq 2\alpha/\sqrt{1+\alpha}$ for functions $f\in\mathcal{S^*}_{\sigma,\alpha}.$
\end{remark}

\begin{definition} A function $f$ given by (\ref{eq1}) is said to be in the class $K_{\sigma}(\varphi)$,
if $f$ and $F$  satisfy the subordinations
\[1+\frac{zf''(z)}{f'(z)}\prec \varphi(z)\;\;{\text {and}}\;\; 1+\frac{wF''(w)}{F'(w)}\prec \varphi(w).\]
\end{definition}
Note that $K_{\sigma}((1+(1-2\beta)z)/(1-z)))=:K_\sigma(\beta)\;\;
    (0\leq\beta<1).$
\begin{theorem}\label{th3} If  $f\in K_{\sigma}(\varphi)$, then
\[|a_2|\leq \min{\set{\sqrt\frac{ B_1^2+B_1+|B_2-B_1|}{6}, \frac{B_1}{2}}}\] and \[
|a_3|\leq \min{\set{\frac{ B_1^2+B_1+|B_2-B_1|}{6}, \frac{B_1(3B_1+2)}{12}}}.\]
\end{theorem}

\begin{proof} Since $f\in K_{\sigma}(\varphi)$,  there are analytic functions $r, s:\mathbb{D}\rightarrow\mathbb{D}$,  with $r(0)=0=s(0)$,  satisfying
\begin{equation}\label{p1.12}1+\frac{zf''(z)}{f'(z)}=\varphi(r(z))\;\text {and}\; 1+\frac{wF''(w)}{F'(w)} =\varphi(s(z)).\end{equation}
Let $p$ and $q$ be defined as in (\ref{pq}),
then it is clear from (\ref{p1.12}), (\ref{p2}) and (\ref{p3}) that
\begin{equation}\label{p4.12}1+\frac{zf''(z)}{f'(z)}=\varphi\left(\frac{p(z)-1}{p(z)+1}\right)\;\;\text {and}\;\; 1+\frac{wF''(w)}{F'(w)}=\varphi\left(\frac{q(z)-1}{q(z)+1}\right).\end{equation}
It follows from (\ref{p4.12}), (\ref{p5}) and (\ref{p6}) that
\begin{equation}\label{eq11.12}
 2a_2=\frac{1}{2}B_1p_1,
\end{equation}
\begin{equation}\label{eq12.12}
 6a_3 =B_1p_1a_2+\frac{1}{2}B_1\left(p_2-\frac{1}{2}p_1^2\right)+\frac{1}{4}B_2p_1^2,
\end{equation}
\begin{equation}\label{eq15.12}
 -2a_2=\frac{1}{2}B_1q_1
\end{equation}
and
\begin{equation}\label{eq16.12}
 6(2a_2^2-a_3)=-B_1q_1a_2+\frac{1}{2}B_1\left(q_2-\frac{1}{2}q_1^2\right)+\frac{1}{4}B_2q_1^2.
\end{equation}
Now (\ref{eq11.12}) and (\ref{eq15.12}) yield
\begin{equation}\label{eq17.12}
 p_1=-q_1
\end{equation}
and
\begin{equation}\label{eq18.12}
 4a_2=\frac{B_1(p_1-q_1)}{2}.
\end{equation}
From (\ref{eq12.12}), (\ref{eq16.12}), (\ref{eq17.12}) and (\ref{eq11.12}), it follows that
\begin{equation}\label{eq19.12}
 48a_2^2=2B_1^2p_1^2+2B_1(p_2+q_2)+(B_2-B_1)(p_1^2+q_1^2).
\end{equation}
In view of $|p_i|\leq2$ and $|q_i|\leq2$ together with (\ref{eq18.12}) and (\ref{eq19.12}) yield the desired estimate on $a_2$ as asserted in the theorem.
In order to find $a_3$, we subtract (\ref{eq12.12}) from (\ref{eq16.12}) and use (\ref{eq17.12}) to obtain
\begin{equation}\label{eq19.22}
 -12a_3=-12a_2^2+\frac{B_1(q_2-p_2)}{2}.
\end{equation}
 Now a computation using (\ref{eq19.12}) and (\ref{eq19.22}) leads to
\begin{equation}\label{eq19.32}
 -48a_3=2B_1^2p_1^2-4B_2p_2+(B_1-B_2)(p_1^2+q_1^2).
\end{equation}
From (\ref{eq18.12}) and (\ref{eq19.22}), it follows that
\begin{equation}\label{eq19.42}
-12a_3=\frac{B_1(q_2-p_2)}{2}-\frac{3(p_1-q_1)^2B_1^2}{16}.
\end{equation}
 Now (\ref{eq19.32}) and (\ref{eq19.42}) yield the desired estimate on $a_3$ as asserted in the theorem.
\end{proof}
\begin{remark}If $f\in K_\sigma(\beta)\;\;(0\leq \beta<1),$ then theorem~\ref{th3} gives
\[|a_2|\leq \min{\set{\sqrt\frac{ (1-\beta)(3-2\beta)}{3}, 1-\beta}}=1-\beta\] and \[
|a_3|\leq \min{\set{\frac{ (1-\beta)(3-2\beta)}{3}, \frac{(1-\beta)(4-3\beta)}{3}}}=\frac{(1-\beta)(3-2\beta)}{3},\] which improves the Brannan and Taha's~\cite[Theorem 4.1]{branan0} estimates $|a_2|\leq \sqrt{1-\beta} $ and $|a_3|\leq 1-\beta$ for functions $f\in K_\sigma(\beta)$.
\end{remark}
\begin{theorem}\label{th12} Let $f\in\sigma$ be given by (\ref{eq1}). If $f\in\mathcal{K}(\varphi)$ and $F\in\mathcal{R}(\varphi)$, then
\[|a_2|\leq \sqrt \frac{3[B_1+|B_2-B_1|]}{8}\] and \[|a_3|\leq\frac{5[B_1+|B_2-B_1|]}{12}.\]
\end{theorem}
\begin{proof} Since $f\in\mathcal{K}(\varphi)$ and $F\in\mathcal{R}(\varphi)$, there exist two analytic functions $r, s:\mathbb{D}\rightarrow\mathbb{D}$,  with $r(0)=0=s(0)$,  such that
\begin{equation}\label{p11.1}1+\frac{zf''(z)}{f'(z)}=\varphi(r(z))\;\text {and}\; F'(w)=\varphi(s(z)).\end{equation}
Let the functions $p$ and $q$ are defined by (\ref{pq}). It is clear that $p$ and $q$ are analytic in $\mathbb{D}$
and $p(0)=1=q(0)$. Also $p$ and $q$  have positive real part in $\mathbb{D}$, and
hence $|p_i|\leq2$ and  $|q_i|\leq2$.
Proceeding as in the proof of Theorem~\ref{th1} it follow from (\ref{p11.1}), (\ref{p5}) and (\ref{p6}) that
$$ 2a_2=\frac{1}{2}B_1p_1,$$
\begin{equation}\label{eq3}
 6a_3-4a_2^2 =\frac{1}{2}B_1\Big(p_2-\frac{1}{2}p_1^2\Big)+\frac{1}{4}B_2p_1^2,
\end{equation}
$$-2 a_2=\frac{1}{2}B_1q_1 $$
and
\begin{equation}\label{eq5}
 3(2a_2^2-a_3)=\frac{1}{2}B_1\Big(q_2-\frac{1}{2}q_1^2\Big)+\frac{1}{4}B_2q_1^2.
\end{equation}
A computation using (\ref{eq3}) and ({\ref{eq5}}), leads to
\begin{equation}\label{eq7}
  a_2^2=\frac{2(p_2+2q_2)B_1+(p_1^2+2q_1^2)(B_2-B_1)}{32}.
\end{equation}
and
\begin{equation}\label{eq8} a_3=\frac{2(3p_2+2q_2)B_1+(3p_1^2+2q_1^2)(B_2-B_1)}{48}.\end{equation}
Now the desired estimates on $a_2$ and $a_3$, follow  from (\ref{eq7}) and (\ref{eq8}) respectively.\end{proof}
\begin{remark} If $f\in\mathcal{K}(\beta)$ and $F\in\mathcal{R}(\beta)$, then from Theorem~\ref{th12} we see that
\[ |a_2|\leq \sqrt {3(1-\beta)}/2\;\;{\text{and}}\;\; |a_3|\leq 5(1-\beta)/6.\]
 In particular if $f\in\mathcal{K}$ and $F\in\mathcal{R}$, then $|a_2|\leq \sqrt 3/2\approx 0.867\;\;{\text{and}}\;\; |a_3|\leq 5/6\approx0.833$.
\end{remark}

\begin{theorem}\label{th21} Let $f\in\sigma$ be given by (\ref{eq1}). If $f\in\mathcal{S}^*(\varphi)$ and $F\in\mathcal{R}(\varphi),$ then
$$|a_2|\leq \frac{\sqrt {5[B_1+|B_2-B_1|]}}{3},\;\;{\text{and} }\;\; |a_3|\leq\frac{7[B_1+|B_2-B_1|]}{9}.$$
\end{theorem}
\begin{proof} Since
   $f\in\mathcal{S}^*(\varphi)$ and $F\in\mathcal{R}(\varphi),$ there exist two analytic functions $r, s:\mathbb{D}\rightarrow\mathbb{D}$,  with $r(0)=0=s(0)$,  such that
\begin{equation}\label{tp1}\frac{zf'(z)}{f(z)}=\varphi(r(z))\;\text {and}\; F'(w)=\varphi(s(z)).\end{equation}
Let the functions $p$ and $q$ be defined as in (\ref{pq}). Then
 {\small\begin{equation}\label{tp4}\frac{zf'(z)}{f(z)}=\varphi\left(\frac{p(z)-1}{p(z)+1}\right)\;\;\text {and}\;\; F'(w)=\varphi\left(\frac{q(w)-1}{q(w)+1}\right).\end{equation}}
 It follow from (\ref{tp4}), (\ref{p5}) and (\ref{p6}) that
  $$a_2=\frac{1}{2}B_1p_1,$$
\begin{equation}\label{teq3}
 2a_3-a_2^2 =\frac{1}{2}B_1\Big(p_2-\frac{1}{2}p_1^2\Big)+\frac{1}{4}B_2p_1^2,
\end{equation}
$$ -2a_2=\frac{1}{2}B_1q_1, $$
\begin{equation}\label{teq5}
 3(2a_2^2-a_3)=\frac{1}{2}B_1\Big(q_2-\frac{1}{2}q_1^2\Big)+\frac{1}{4}B_2q_1^2.
\end{equation}
A computation using (\ref{teq3}) and ({\ref{teq5}}) leads to
\begin{equation}\label{teq7}
  a_2^2=\frac{2(3p_2+2q_2)B_1+(3p_1^2+2q_1^2)(B_2-B_1)}{36}
\end{equation} and
\begin{equation}\label{teq8} a_3=\frac{2(6p_2+q_2)B_1+(6p_1^2+q_1^2)(B_2-B_1)}{36}.\end{equation}
Now the bounds for $a_2$ and $a_3$ are obtained from (\ref{teq7}) and (\ref{teq8}) respectively using the fact that $|p_i|\leq 2$ and $|q_i|\leq 2$.
\end{proof}
\begin{remark}
 If $f\in \mathcal{S}^*(\beta)$ and $F\in\mathcal{R}(\beta)$, then from Theorem~\ref{th21} it is easy to see that \[ |a_2|\leq \sqrt{10(1-\beta)}/3 \;\;{\text{and}}\;\;|a_3|\leq 14(1-\beta)/9.\] In particular if $f\in \mathcal{S}^*$ and $F\in\mathcal{R}$, then $|a_2|\leq \sqrt 10/3\approx1.054 \;\;{\text{and}}\;\;|a_3|\leq14/9\approx 1.56.$
\end{remark}
\begin{theorem}\label{th5} Let $f\in\sigma$ given by (\ref{eq1}). If $f\in\mathcal{S}^*(\varphi)$ and $F\in\mathcal{K}(\varphi),$ then
\[|a_2|\leq \sqrt \frac{B_1+|B_2-B_1|}{2}\] and
 \[|a_3|\leq\frac{B_1+|B_2-B_1|}{2}.\]
\end{theorem}
\begin{proof}
  Assuming $f\in\mathcal{S}^*(\varphi)$ and $F\in\mathcal{K}(\varphi)$ and proceeding in the similar way as in the proof of Theorem~\ref{th12}, it is easy to see that
  $$ a_2=\frac{1}{2}B_1p_1,$$
\begin{equation}\label{te2}
 3a_3-a_2^2 =\frac{1}{2}B_1\Big(p_2-\frac{1}{2}p_1^2\Big)+\frac{1}{4}B_2p_1^2,
\end{equation}
$$ -2a_2=\frac{1}{2}B_1q_1,$$
\begin{equation}\label{te4}
 8a_2^2-6a_3=\frac{1}{2}B_1\Big(q_2-\frac{1}{2}q_1^2\Big)+\frac{1}{4}B_2q_1^2.
\end{equation}
A computation using (\ref{te2}) and (\ref{te4}) leads to
\begin{equation}\label{te7}
 a_2^2=\frac{2(2p_2+q_2)B_1+(2p_1^2+q_1^2)(B_2-B_1)}{24}
\end{equation} and
\begin{equation}\label{te8}
 a_3=\frac{2(8p_2+q_2)B_1+(8p_1^2+q_1^2)(B_2-B_1)}{72}.
\end{equation}
Now using the result $|p_i|\leq2$ and $|q_i|\leq2$, the estimates on $a_2$ and $a_3$ follow from (\ref{te7}) and (\ref{te8}) respectively.
\end{proof}
\begin{remark} Let $f\in\mathcal{S}^*(\beta)$ and $F\in\mathcal{K}(\beta),\;0\leq\beta<1$. Then from Theorem~\ref{th5}, it is easy to see that \[ |a_2|\leq \sqrt {1-\beta}\;\; {\text{and}}\;\; |a_3|\leq 1-\beta.\] In particular if $f\in\mathcal{S}^*$ and $F\in\mathcal{K},$ then $|a_2|\leq 1\;\; {\text{and}}\;\; |a_3|\leq 1$.
\end{remark}

\noindent{\bf Acknowledgements.} The research is supported by a grant from University of Delhi.

\end{document}